OpenStudy (kmallette09):

The width and length of a rectangle (in feet)are consecutive odd integers. If the length is increased by 5 feet, the area of the resulting rectangle is 60 square feet. What is the area of the original rectangle? 25 ft 2 30 ft 2 35 ft 2

OpenStudy (kmallette09):

ft 2 means squared

OpenStudy (will.h):

ii would suggest you to do the following. 1st find the are of all the options Area= length * width 1st option= 25* 2 1st option= 50 (makes sense because it's lower than 60) 2nd option= 30*2 2nd option= 60 (also make sense) 3rd option= 35*2 3rd option= 70 (doesn't make sense because if we subtract 5 from the length it would be 30 *2 which equals 60 and since the original rectangle are cannot bee the same ad the new one then you can safely scratch this option) Now it's between 1 and 2

OpenStudy (kmallette09):

so is it 35ft squared

OpenStudy (will.h):

if we subtract 5 from the length of the 1st option we would get 20*2 which equals 40 which is a possible answer and if we also subtract 5 from the 2nd option we would get 25*2 which equals 50 So that's why am Confused both 1 and 2 can work

OpenStudy (kmallette09):

i think its 25ft squared then

OpenStudy (will.h):

since it's the closest to 60ft then, i guess you right :)

OpenStudy (kmallette09):

can u help be on the next question 2

OpenStudy (will.h):

Sure post them and tag me :)

OpenStudy (kmallette09):

do u want me to close this question

OpenStudy (will.h):

since we solved it, then i guess it's better to.

OpenStudy (will.h):

i'll be waiting for your other questions tho.

OpenStudy (kmallette09):

k